fourteen is three times the difference of a number and 6. write an equations
Answer by stanbon(75887)

fourteen is three times the difference of a number and 6. write an equations
-----------
3(x-6) = 14
3x-18 = 14
3x = 32
x = 32/3 or 10 2/3
================
Cheers,
Stan H.